OpenStudy (anonymous):

Solve the system of equations by substitution. What is the solution for x? 2x + y = 1 4x + 2y = -2 x = 0 x = 2 There is no x value as there is no solution. x can be any value as there is an infinite number of solutions.

OpenStudy (anonymous):

@ganeshie8

OpenStudy (tkhunny):

Well, do it. Solve the top equation for y, since that looks easiest and substitute away!

OpenStudy (nirmalnema):

2x+y=1 4x+2y=-2 = 2x+y=-1 these are 2 parallel line since it has no solution

OpenStudy (tkhunny):

@nirmalnema It's a lovely result, but where is your substitution?

OpenStudy (nirmalnema):

you will not get any ans ....by substitution

OpenStudy (tkhunny):

That is not correct. The answer is unique. Substitution WILL get you there. OP doesn't seem interested, though.

OpenStudy (tkhunny):

2x + y = 1 4x + 2y = -2 Solving the 1st y = 1 - 2x Substituting into the second 4x + 2(1-2x) = -2 Simplify 2x + (1-2x) = -1 2x + 1 - 2x = -1 2x - 2x = -1-1 0 = -2 As this is entirely false, we must conclude that there is no solution to this system. It is a matter of interpreting the result correctly.